Grade 12Integral Calculus

∫lnx dx

Lovey
12 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int lnxdx
I = \int x^{0}lnxdx
Apply ILATE Rule
I = xlnx - \int dx
I = xlnx - x + c
I = x(lnx - 1) + c